given that log[a]5=0.8271 and log[a]20=1.5395 fing log[a]4
Answer by mickclns(59)

The important formula here is log[d](b/c) = log[d]b - log[a]c, so letting d = a, b=20 and c = 5,
log[a]4 = log[a](20/5) = log[a]20 - log[a]5 = 1.5395 - 0.8271 = 0.7124